# Improving the Reasoning of Multi-Image Grounding in MLLMs via Reinforcement Learning

### v1

@@ Abstract
Recently, Multimodal Large Language Models (MLLMs) excel at visual grounding in single-image scenarios with textual references. However, their performance degrades when handling real-world applications involving complex multi-image compositions and multimodal instructions, which reveals limitations in cross-image reasoning and generalization. To address these challenges, we adopt a Reinforcement Learning (RL) based post-training strategy to improve the reasoning performance of MLLMs in multi-image grounding tasks. Our approach begins with synthesizing high-quality chain-of-thought (CoT) data for cold-start initialization, followed by supervised fine-tuning (SFT) using low-rank adaptation (LoRA). The cold-start training stage enables the model to identify correct solutions. Subsequently, we perform rejection sampling using the merged SFT model to curate high-quality RL data and leverage rule-based RL to guide the model toward optimal reasoning paths. Extensive experimental results demonstrate the effectiveness of our approach, achieving +9.04% improvements on MIG-Bench and +4.98% improvements on several out-of-domain reasoning grounding benchmarks over the SFT baseline. Furthermore, our approach exhibits strong generalization in multi-image perception, with gains of +3.1% and +2.4% over the base model on subsets of the BLINK and MMIU benchmarks, respectively.
@@ 1 Introduction
Figure 1: Top: Examples of our multi-image reasoning grounding. Bottom: The comparison of Qwen2.5-VL-7B model, SFT model, and our reasoning model. Our method achieves the best performance among all benchmarks.
The traditional visual grounding task focuses on localizing target regions within a single image using simple natural language descriptions. Recent research leveraging large language models (LLMs) has significantly improved performance on this task by utilizing LLMs’ powerful language comprehension capabilities. However, real-world applications often require sophisticated multi-image grounding with complex instructions, including cross-image reasoning. As illustrated in Figure 1, practical scenarios often require identifying the region that best matches the semantics of the query image. Such tasks demand advanced multi-modal reasoning and comprehensive understanding across multiple visual contexts.
Recently, Migician li2025migician introduces a multi-image grounding dataset with the MIG-Bench benchmark. By employing a two-stage supervised fine-tuning (SFT) pipeline on their dataset, they have achieved impressive results, demonstrating the potential of MLLMs for multi-image grounding tasks. However, their SFT-based approach primarily “memorize” chu2025sft understanding patterns and instruction-following behaviors in multi-image scenarios, rather than developing robust reasoning capabilities. This limitation constrains the model’s ability to generalize to novel scenarios and real-world applications.
@@
Inspired by the recent success in RL-based post-training frameworks for Large Reasoning Models (LRM)  guo2025deepseek ; team2025kimi ; yang2025qwen3 , we explore the potential of the RL paradigm to improve the reasoning abilities of MLLMs in multi-image grounding scenarios. Our preliminary attempts applied RL directly to the state-of-the-art open-source MLLM, Qwen2.5-VL-7B bai2025qwen2 . However, we found that the model often failed to generate correct responses during the RL training stage, which we attribute to two key limitations of the base model: (1) inadequate capacity for processing complex image queries, and (2) insufficient ability to handle multi-image contexts. It suggests that directly employing RL paradigm on the base model is ineffective for exploring reasoning paths in such challenging scenarios.
@@
To overcome these limitations, we synthesize high-quality chain-of-thought (CoT) data for cold-start initialization using large-scale MLLMs, Qwen2.5-VL-72B. This data serves for supervised fine-tuning (SFT) with low-rank adaptation (LoRA), enabling the model to develop robust multi-image comprehension and reasoning capabilities. Then, we remove low-quality data through rejection sampling according to predictions from the merged cold-start SFT model. Finally, we employ a rule-based RL paradigm based on Group Relative Policy Optimization (GRPO) shao2024deepseekmath to guide the model toward discovering correct reasoning paths while improving generalization performance.
@@
We conduct comprehensive experiments on several benchmarks to demonstrate the effectiveness of our training pipeline. On the multi-image grounding benchmark MIG-Bench li2025migician , our model achieves improvements of +3.46% (in-domain) and +12.76% (out-of-domain) over the SFT baseline. For zero-shot grounding evaluation, our approach outperforms the SFT baseline by +6.23% on Ref-L4 chen2025revisiting , +4.79% on LISA-Grounding lai2024lisa , and +2.83% on ReVOS-Grounding yan2024visa . Furthermore, for the multi-image understanding benchmarks, our approach exhibits strong generalization capabilities, with improvements of +3.1% on BLINK fu2024blink and +2.4% on MMIU meng2024mmiu compared to the base Qwen2.5-VL-7B model.
@@ 2 Related Works
Referring grounding, also known as referring expression comprehension (REC), involves identifying a specific object within an image based on a given referring natural language expression. With the rapid advancement of multimodal large language models (MLLMs) liu2023visual ; wang2024qwen2 ; li2024llava ; jiang2024mantis ; li2024llavainterleave ; bai2025qwen2 ; zhu2025internvl3 , significant progress has been made in extending their capabilities to address more challenging tasks, including fine-grained image understanding, object detection zhan2024griffon ; jiao2024lumen , and visual grounding lai2024lisa ; yan2024visa ; rasheed2024glamm ; li2024groundinggpt . Although these MLLMs have achieved strong performance on standard benchmarks such as RefCOCO/+/g mao2016generation ; yu2016modeling , their visual grounding capabilities remain largely confined to single-image contexts and face substantial difficulties in capturing cross-image relationships. Recent efforts, such as Migician li2025migician , introduce multi-image grounding tasks and further propose a two-stage supervised fine-tuning strategy, achieving notable results in these tasks. However, such approaches still exhibit insufficient reasoning abilities and limited generalization in multi-image and out-of-domain settings. In this paper, we explore the potential of the RL paradigm to improve the reasoning and comprehension abilities of MLLMs in multi-image grounding scenarios.
2.2 Reinforcement Learning-based Reasoning Model
@@
Recent advancements in improving the reasoning ability of large language models jaech2024openai ; xu2024llava ; el2025competitive ; guo2025deepseek ; team2025kimi have been made through Chain-of-Thought (CoT) training or reinforcement learning-based post-training. Previous methods, such as Reinforcement Learning from Human Feedback (RLHF) ouyang2022training and Direct Preference Optimization (DPO) rafailov2023direct , have explored aligning models with human preferences, achieving remarkable success in enhancing reasoning and instruction-following capabilities. Recently, DeepSeek-R1 guo2025deepseek employed the Group Relative Policy Optimization (GRPO) shao2024deepseekmath algorithm as a post-training method without the need for critic models, demonstrating the potential of large-scale reinforcement learning. Inspired by the success of DeepSeek-R1, a series of subsequent works rapidly adopted this approach to enhance multi-modal reasoning across various domains, including mathematical problem solving yang2025r1 ; peng2025skywork ; huang2025vision ; zhang2025r1 ; deng2025openvlthinker , video understanding feng2025video ; zhang2025tinyllava ; liao2025improved ; wang2025time , and visual perception tasks shen2025vlm ; liu2025seg ; deng2025boosting ; liu2025visual ; zhan2025vision ; yu2025perception ; tan2025reason . Among the most closely related works to ours, VLM-R1 shen2025vlm , Visual-RFT liu2025visual , Vision-R1 zhan2025vision , Perception-R1 yu2025perception , most focus exclusively on single-image visual grounding or object detection. However, these approaches lack the capability for multi-image perception and reasoning. In this paper, we follow the DeepSeek-R1 paradigm and extend it to the multi-image grounding tasks, while also achieving strong performance on single-image grounding and multi-image understanding tasks.
@@ 3 Method
3.1 Basic Settings
We begin by formally defining the multi-image grounding task. Formally, given a natural language description $t$, a query image $Q$, and several target images $\{I_{1},I_{2},...,I_{m}\}$, the model M is required to generate a bounding box $G$ that indicates the region satisfying the semantic and contextual constraints specified by $t$ and $Q$.
Unlike conventional grounding tasks that rely on fixed instruction templates, multi-image grounding involves dynamic and flexible instructions, which demand significantly stronger instruction comprehension and reasoning capabilities from MLLMs. To address the challenge, we introduce a reinforcement learning training strategy to enhance the model’s reasoning capabilities, inspired by recent breakthrough improvements in LRMs. We employ Qwen2.5-VL-7B as the base model due to its excellent multi-modal comprehension performance. Our RL paradigm is similar to DeepSeek-R1 guo2025deepseek , consisting of cold-start CoT-SFT initialization, rejection sampling, and rule-based RL training. We will provide detailed descriptions of our dataset construction strategy and training paradigm in the following subsections.
3.2 Cold Start Initialization
As analyzed in DeepSeek-R1 guo2025deepseek , directly applying reinforcement learning (RL) from scratch may cause the model to enter an unstable state. Specifically, the base MLLM model struggles to handle complex multi-image contexts and accurately localize target regions within multi-image scenarios. To mitigate these issues, it is crucial to pre-equip the model with task-specific reasoning capabilities before RL fine-tuning. We achieve this by constructing a high-quality CoT cold-start dataset, which provides structured reasoning guidance to bootstrap the model’s initial performance.
@@ 3.2 Cold Start Initialization
3.2.1 Chain-of-Thought Reasoning Data Collection
We construct our CoT dataset primarily based on MGrounding-630k dataset li2025migician . To maintain the model’s capability in single-image grounding scenarios, additionally introduce a small amount of single-image grounding data, including RefCOCO/+/g mao2016generation ; yu2016modeling , as well as data from other sources, such as ODINW li2022grounded and V3Det wang2023v3det .
We utilize the advanced large-scale MLLM Qwen2.5-VL-72B to generate CoT reasoning data. Given a question, along with images and a fixed prompt similar to DeepSeek-R1, the model generates CoT rationales processes in the following format:
To improve data quality, we conduct an additional data filtering strategy. For each sample, we prompt the model to produce four distinct responses. Then, We estimate the difficulty of each sample based on the Intersection over Union (IoU) accuracy among these four responses, retaining only samples for which the accuracy is 1 to ensure the reliability of the data. After filtering, we obtain 56k high-reliability CoT cold-start samples, with detailed statistics presented in Table 1.
3.2.2 Supervised Fine-tuning
Leveraging our CoT cold-start dataset, we perform fine-tuning of the Qwen2.5-VL model to develop its multi-image comprehension and reasoning capabilities. During this training phase, we implement strong supervision via next-token prediction across the entire generation process, effectively incorporating knowledge at both the reasoning trace level and the final bounding box prediction.
We recognize that conventional full-parameter fine-tuning risks locking the model into rigid reasoning patterns, which could compromise both its generalization capacity and conversational fluency. To address this, we employ the LoRA fine-tuning approach, which maintains the model’s foundational knowledge while significantly enhancing its perception and reasoning performance in multi-image grounding tasks. Then we merge the LoRA parameters with the original base model, yielding our stage-1 CoT-SFT model.
3.3 Reinforcement Learning
In the second stage, we implement GRPO shao2024deepseekmath , a rule-based reinforcement learning training paradigm, to further enhance the model’s reasoning capabilities. The model samples multiple candidate responses for each sample and are encouraged to reinforce responses achieving higher task-level rewards during GRPO training stage. To be specific, our model generates a group of $G$ complete responses $\{o_{1},o_{2},...,o_{G}\}$ with current parameter group $\pi_{\theta}$. Each response contains a full CoT reasoning process and a final bounding box prediction. For each $o_{i}$, we compute a scalar reward $r_{i}$, and normalize these rewards to estimate group-relative advantages $A_{i}$, formally as:
@@ 3.3 Reinforcement Learning
Then, the training objective is given as follows:
@@
where $\pi_{\theta_{old}}$ and $\pi_{ref}$ are the parameter groups of old model without latest updating and the fixed stage-1 CoT-SFT model. The parameter $\epsilon$ is used to control the clipping range, and $\beta$ is the KL penalty coefficient.
@@
To ensure efficient training, we implement a rejection sampling strategy based on our stage-1 CoT-SFT model. Specifically, we perform multiple sampling on the RL training data and remove samples where the model’s predictions are all correct or wrong. This approach is adopted because samples that the model answers either completely incorrectly or completely correctly do not contribute effectively to the relative advantage estimation in the GRPO algorithm. Retaining only samples where the model provides partially correct answers enables more efficient model optimization. After rejection sampling, we collect 174k data for stage-2 training.
To optimize the training efficiency of our reinforcement learning framework, we employ a rejection sampling strategy using our stage-1 CoT-SFT model. Specifically, we generate multiple predictions for each sample in the RL training dataset. Those samples where all predictions are correct or wrong are removed, as they contribute poorly to the relative advantage estimation in GRPO. On the other hand, partially correct samples offer optimal learning signals for model improvement. After applying this rigorous filtering process, we obtain a high-quality dataset of 174,000 samples for stage-2 RL training.
Reward Function
@@
The total reward is computed as a weighted sum of the accuracy and format rewards:

### v2

@@ Abstract
Recently, Multimodal Large Language Models (MLLMs) excel at visual grounding in single-image scenarios with textual references. However, their performance degrades when handling real-world applications that involve complex multi-image compositions and multi-modal instructions, revealing limitations in cross-image reasoning and generalization. To address these challenges, we adopt a Reinforcement Learning (RL) based post-training strategy to improve the reasoning of MLLMs in multi-image grounding tasks. Our approach begins with synthesizing high-quality chain-of-thought (CoT) data for cold-start initialization, followed by supervised fine-tuning (SFT) using low-rank adaptation (LoRA). The cold-start training stage enables the model to identify correct solutions. Subsequently, we perform rejection sampling using the merged SFT model to curate high-quality RL data and leverage rule-based RL to guide the model toward optimal reasoning paths. Extensive experimental results demonstrate the effectiveness of our approach, yielding improvements of +9.04% on MIG-Bench, +6.37% on MC-Bench, and +4.98% on several out-of-domain reasoning grounding benchmarks compared to the SFT baseline. Furthermore, our method exhibits strong generalization in multi-image perception, with gains of +3.1% and +2.4% over the base model on BLINK and MMIU benchmarks, respectively.
The traditional visual grounding task focuses on localizing target regions within a single image using simple natural language descriptions. Recent research leveraging large language models (LLMs) has significantly improved performance on this task through their powerful language comprehension capabilities. However, real-world applications often require sophisticated multi-image grounding with complex instructions, including cross-image reasoning. As illustrated in Figure 1 (a), practical scenarios often require identifying the region that best matches the semantics of the query image. Such tasks demand advanced multi-modal reasoning and comprehensive multi-image understanding.
Recently, Migician li2025migician introduces a multi-image grounding dataset and employs a two-stage supervised fine-tuning (SFT) pipeline on their dataset. However, their SFT approach primarily “memorize” chu2025sft understanding patterns and instruction-following behaviors in multi-image scenarios, rather than developing reasoning capabilities. This limitation constrains the model’s generalization to novel scenarios and real-world applications.
@@ 1 Introduction
Figure 1: (a): Example of our multi-image reasoning grounding results. (b): The comparison of the Qwen2.5-VL-7B model, its SFT variant, and our reasoning model. Our method achieves the best performance among all benchmarks.
@@
Inspired by the recent success in RL-based post-training frameworks for Large Reasoning Models (LRM)  guo2025deepseek ; team2025kimi ; yang2025qwen3 , we explore the potential of the RL paradigm to improve MLLMs’ reasoning abilities in multi-image grounding scenarios. Our preliminary attempts applied RL directly to the superior open-source MLLM, Qwen2.5-VL-7B bai2025qwen2 . However, we found that the model often failed to generate correct responses during RL training, due to the limited capacity of the base model to process complex image queries and multi-image contexts.
@@
To address these limitations, we introduce a two-stage training framework, consisting of cold-start CoT-SFT initialization and rule-based reinforcement learning. Firstly, we synthesize high-quality chain-of-thought (CoT) data using a large-scale MLLM, Qwen2.5-VL-72B, and perform supervised fine-tuning with LoRA for cold-start initialization, enabling the model to acquire basic multi-image reasoning capabilities. Next, we adopt a rule-based RL paradigm built on Group Relative Policy Optimization (GRPO) shao2024deepseekmath to guide the model toward discovering correct reasoning paths and improving generalization. However, we observe that a significant portion of data is low-quality, preventing the model from generating response candidates with sufficient relative advantage gaps, thereby providing limited supervisory signals for RL. To improve training efficiency, we apply rejection sampling, according to predictions from the merged cold-start SFT model, to filter out such uninformative samples before the RL stage.
@@
We conduct comprehensive experiments on several benchmarks to demonstrate the effectiveness of our training pipeline. As shown in Figure 1 (b), our model achieves improvements of +3.46% (in-domain) and +12.76% (out-of-domain) over the SFT baseline on the multi-image grounding benchmark MIG-Bench li2025migician . As for the other multi-image grounding benchmark MC-Bench xu2024mc , our model demonstrates a +6.37% performance gain compared to the SFT baseline. For zero-shot grounding evaluation, our approach outperforms the SFT baseline by +4.79% on LISA-Grounding lai2024lisa , +6.23% on Ref-L4 chen2025revisiting , and +2.83% on ReVOS-Grounding yan2024visa . Furthermore, for the multi-image understanding benchmarks, our approach demonstrates strong generalization capabilities, with improvements of +3.1% on BLINK fu2024blink and +2.4% on MMIU meng2024mmiu compared to the base Qwen2.5-VL-7B model.
@@ 2 Related Works
Referring grounding, also known as referring expression comprehension (REC), involves identifying a specific object within an image based on a given natural language referring expression. With the rapid advancement of multimodal large language models (MLLMs) liu2023visual ; wang2024qwen2 ; li2024llava ; jiang2024mantis ; li2024llavainterleave ; bai2025qwen2 ; zhu2025internvl3 , significant progress has been made in extending their capabilities to address more challenging tasks, including fine-grained image understanding wang2024advancing , object detection zhan2024griffon ; jiao2024lumen , and visual grounding lai2024lisa ; yan2024visa ; rasheed2024glamm ; li2024groundinggpt . Although these MLLMs have achieved strong performance on standard benchmarks such as RefCOCO/+/g mao2016generation ; yu2016modeling , their visual grounding capabilities remain largely confined to single-image contexts and face substantial difficulties in capturing cross-image relationships. Recent efforts, such as Migician li2025migician , introduce multi-image grounding tasks and further propose a two-stage SFT strategy, achieving notable results in these tasks. However, such approaches still exhibit insufficient reasoning abilities and limited generalization in multi-image and out-of-domain settings. In this paper, we explore the potential of the RL paradigm to improve the reasoning and comprehension abilities of MLLMs in multi-image grounding scenarios.
2.2 RL-based Reasoning Model
@@
Recent advancements in improving the reasoning ability of large language models jaech2024openai ; xu2024llava ; el2025competitive ; guo2025deepseek ; team2025kimi have been made through CoT-based or RL-based post-training. Previous methods, such as Reinforcement Learning from Human Feedback (RLHF) ouyang2022training and Direct Preference Optimization (DPO) rafailov2023direct , have explored aligning models with human preferences, achieving remarkable success in enhancing reasoning and instruction-following capabilities. Recently, DeepSeek-R1 guo2025deepseek employes the GRPO shao2024deepseekmath algorithm as a post-training method without the need for critic models, demonstrating the potential of large-scale reinforcement learning. Inspired by the success of DeepSeek-R1, a series of subsequent works rapidly adopt this approach to enhance multi-modal reasoning across various domains, including mathematical problem solving yang2025r1 ; peng2025skywork ; huang2025vision ; zhang2025r1 ; deng2025openvlthinker , video understanding feng2025video ; zhang2025tinyllava ; liao2025improved ; wang2025time , and visual perception tasks shen2025vlm ; liu2025seg ; deng2025boosting ; liu2025visual ; zhan2025vision ; yu2025perception ; tan2025reason . Among the most closely related works to ours, VLM-R1 shen2025vlm , Visual-RFT liu2025visual , Vision-R1 zhan2025vision , and Perception-R1 yu2025perception , primarily focus on single-image visual grounding or object detection. However, these approaches lack the capability for multi-image perception and reasoning. In this paper, we follow the DeepSeek-R1 paradigm and extend it to the multi-image grounding tasks, while also achieving strong performance on both multi-image grounding and multi-image understanding tasks.
@@ 3 Method
3.1 Preliminaries
We begin by formally defining the multi-image grounding task. Given a natural language description $t$, query images $Q$, and a set of $m$ target images $\{I_{1},I_{2},...,I_{m}\}$, the model M is required to localize a region of interest by generating a bounding box $G$ that satisfying the semantic and contextual constraints specified by $t$ and $Q$.
While traditional grounding tasks typically rely on fixed instruction templates, the multi-image grounding task introduces the additional complexity of processing dynamic, context-sensitive instructions. This advancement necessitates MLLMs with substantially enhanced comprehension and reasoning abilities. To bridge this capability gap, we adopt an RL-based training paradigm, leveraging recent advances in LRMs, which have demonstrated significant improvements in complex reasoning tasks. Specifically, we choose Qwen2.5-VL-7B as the base model due to its excellent multi-modal comprehension performance. As shown in Figure 2, our training paradigm consists of two stages: cold-start CoT-SFT initialization and rule-based RL training. We will provide detailed descriptions of each stage in the following subsections.
3.2 Cold-start CoT-SFT Initialization
As analyzed in DeepSeek-R1 guo2025deepseek , directly applying reinforcement learning from scratch may result in unstable convergence. Specifically, the base model struggles to handle complex multi-image contexts and accurately localize target regions in multi-image scenarios. Consequently, it is essential to pre-equip the MLLM with task-specific reasoning capabilities before RL training. To address this, we construct a high-quality CoT cold-start dataset that provides reasoning guidance to bootstrap the model’s initial performance.
Figure 2: The overview of our two-stage training paradigm. Our training paradigm consists of a cold-start CoT-SFT initialization (stage-1) and a rule-based RL training (stage-2).
CoT Reasoning Dataset
Our CoT reasoning dataset is constructed based on MGrounding-630k dataset li2025migician . To maintain the model’s capability in single-image grounding scenarios, we augment our dataset with a small amount of single-image grounding data RefCOCO/+/g mao2016generation ; yu2016modeling , as well as object detection datasets: ODINW li2022grounded and V3Det wang2023v3det . This hybrid approach enables comprehensive training across both multi-image and single-image grounding tasks.
We employ a large-scale MLLM Qwen2.5-VL-72B to generate the CoT reasoning data. Following the paradigm of DeepSeek-R1, our generation process takes a question-image pair accompanied by a fixed instructional prompt as input. The MLLM then produces structured CoT rationales in the following format:
To establish high-quality training data, we implement an additional filtering strategy. For each sample, we prompt the Qwen2.5-VL-72B model to generate four distinct responses. Next, we assess sample consistency by computing the Intersection over Union (IoU) accuracy across all responses. To guarantee maximum reliability in our training set, we retain only samples achieving perfect agreement (IoU accuracy = 1.0). This filtering process yields 56k high-confidence CoT cold-start samples, with detailed statistics provided in Table 1.
@@ 3.2 Cold-start CoT-SFT Initialization
Cold-start Supervised Fine-tuning
Leveraging our curated CoT dataset, we equip the Qwen2.5-VL-7B model with multi-image comprehension and reasoning capabilities through cold-start initialization. To be specific, we perform strong supervision via next-token prediction across the entire generation process. This stage ensures comprehensive learning of both the reasoning trace and its spatial grounding predictions.
We recognize that full-parameter fine-tuning may constrain the model’s reasoning flexibility, potentially diminishing both its generalization capabilities and the quality of its natural language interactions. To mitigate these limitations, we employ a LoRA fine-tuning approach, which selectively updates only low-rank decomposed matrices. This approach maintains the model’s foundational knowledge while significantly enhancing its perception and reasoning performance in multi-image grounding tasks. After fine-tuning, we merge the learned LoRA modules with the base MLLM, resulting in our Stage-1 CoT-SFT model with improved capabilities.
3.3 Rule-based RL Training
In the second stage, we implement group relative policy optimization (GRPO) shao2024deepseekmath , a rule-based reinforcement learning algorithm designed to improve policy optimization by incorporating group-wise relative constraints on policy updates. To improve training efficiency, we integrate rejection sampling before GRPO, filtering out uninformative samples to accelerate convergence and stabilize learning.
Group Relative Policy Optimization
The MLLM model generates a group of $G$ complete responses $\{o_{1},o_{2},...,o_{G}\}$ with current policy $\pi_{\theta}$. The output response is required to contain a full CoT reasoning process and a final bounding box prediction. For each response $o_{i}$, we compute a scalar reward $r_{i}$, and normalize these rewards to estimate its group-relative advantage $A_{i}$, formally as:
@@ 3.3 Rule-based RL Training
Then, the GRPO training objective can be defined as follows:
@@
where $\pi_{\theta_{old}}$ is the previous policy before update, and $\pi_{ref}$ is the fixed CoT-SFT policy after stage-1 training. The hyperparameters $\epsilon$ and $\beta$ control the policy update clipping range and KL divergence penalty strength, respectively.
Models Referential Grounding Spontaneous Grounding AVG
Visual Reference Textual Visual+Textual Difference Similarity
OT MV Region Refer GG Reason Co-Re Static Robust Common
Human Performance
Human 100.00 96.88 100.00* 98.99 91.06* 92.08 97.44 99.50* 97.87 98.00* 97.18
70B-Scale MLLMs
LLaVA-OV-72B 12.91 7.64 2.14 17.83 21.60 11.88 8.55 13.26 5.34 26.84 13.65
InternVL2-76B 30.73 20.83 5.74 46.46 41.28 32.67 26.50 15.91 10.64 36.40 26.72
InternVL3-78B 27.08 14.58 10.44 50.51 38.08 45.54 17.09 10.04 9.57 24.12 24.71
Qwen2-VL-72B 26.73 22.57 18.62 33.33 62.53 50.50 17.09 46.12 46.81 64.46 38.88
Qwen2.5-VL-72B 34.32 29.17 8.31 62.63 59.92 66.34 41.03 43.75 46.81 69.98 46.23
7B-Scale MLLMs
Mantis 12.18 2.08 1.00 1.01 10.02 0.00 0.85 1.52 0.00 3.31 3.20
LLaVA-OV-7B 0.18 1.04 1.08 9.09 15.43 6.93 0.85 6.06 3.19 3.43 4.73
Minicpm2.6 9.82 6.25 1.75 11.11 10.02 2.97 2.56 14.58 2.13 14.34 7.55
mPLUG-Owl3 8.55 7.64 2.41 7.07 22.85 9.09 5.98 18.56 6.38 34.93 12.35
InternVL2-8B 20.73 9.72 3.49 28.28 30.26 17.82 9.40 6.92 7.45 25.49 15.96
InternVL3-8B 14.84 6.94 12.13 7.07 34.87 16.83 2.56 23.67 14.89 47.99 18.18
Qwen2-VL-7B 20.73 11.81 25.95 23.23 58.52 48.51 11.97 27.84 38.30 19.36 28.62
Qwen2.5-VL-7B 15.23 5.56 4.07 13.13 32.26 3.96 2.56 28.03 5.32 21.81 13.19
+ SFT 26.59 35.07 39.32 79.80 59.72 53.47 23.93 53.22 44.68 82.11 49.79
+ CoT-SFT 46.14 34.03 28.93 80.81 63.93 66.34 34.19 47.35 45.74 76.59 52.41
Ours 63.41 38.54 51.45 82.83 69.54 67.33 37.61 51.52 43.62 82.48 58.83
@@
We implement rejection sampling leveraging our stage-1 CoT-SFT model to enhance the quality of training data. This selective approach involves: (1) generating multiple predictions for each sample in the RL training dataset, and (2) systematically eliminating samples where the model demonstrates either complete correctness (all predictions correct) or failure (all predictions incorrect). The rejection sampling is theoretically grounded in the requirements of the GRPO algorithm, where samples with uniformly correct or incorrect predictions provide negligible gradient information for relative advantage estimation. By retaining only samples with partial correctness, we ensure the training data contains optimal learning signals that effectively drive model optimization. After filtering, we curate a high-quality dataset comprising precisely 174k samples for stage-2 rule-based RL training.
Reward Design
@@
The total reward is computed as a weighted sum of both the accuracy reward and the format reward: